OpenStudy (anonymous):

x-5y=45 the value of y when x equals 5

OpenStudy (anonymous):

the answer is -8

OpenStudy (anonymous):

What value of y would you need when x=5 to keep that statement true? x-5y=45 x=45+5y x-45=5y (5)-45=5y -40=5y y=-8

OpenStudy (ddcamp):

First, solve for y: \[x-5y = 45 \rightarrow -5y = 45 - x \rightarrow y = - \frac{ 45 - x }{ 5 }\] And plug in x: \[- \frac{ 45 - 5 }{ 5 } \rightarrow - \frac{ 40 }{ 5 } \rightarrow -8\]

OpenStudy (anonymous):

thanks